An assembly line with 17 tasks is to be balanced. The longest task is 2.3 minutes, and the total time for all tasks is 18 minutes. The line will operate for 490 minutes per day.

a. What is the minimum number of workstations needed if the output rate is to be 170 units per day? (Round the final answer to the next whole number.)

Number of workstation            

b. What cycle time will provide an output rate of 126 units per day? (Round the final answer to 2 decimal places.)

Cycle time              minutes per unit

c. What output will result if the cycle time is (i) 10 minutes? (ii) 15 minutes? (Round the final answers to the next whole number.)

  Output of 10 Min CT   units per day  
  Output of 15 Min CT   units per day  
Answer #1

a)

Cycle Time for 170 units = Total Available Time/Demand = 490/170 minutes

Theoretical Number of Workstation = Total Task Time/Desired Cycle Time

Theoretical Number of Workstation = 18/(490/170)

Theoretical Number of Workstation = 6.24 = 7 workstations

b)

Cycle Time = Total Available Time/Desired Output = 490/126 = 3.89 minutes

c)

Cycle Time = 10 minutes

Output = Available Time/cycle Time = 490/10 = 49 units

Cycle Time = 15 minutes

Output = 490/15 = 32.67 = 33 units